PuzzledGurl:

Lydia bought a shirt at 20% off its retail price of $40. She paid 5% tax on the price after the discount. How much did Lydia pay for the shirt?

Ultrilliam:

@dude

dude:

Cool Do you know how to get 20% off of $40?

PuzzledGurl:

Uhm the only way I know is to look it up... so in other words no

dude:

Okay, so first you want to convert the 20% to a decimal, do you know how to do that?

PuzzledGurl:

0.20

PuzzledGurl:

0.20

dude:

That looks good, now to know what 20% of the $40 is, multiply the decimal and the price

PuzzledGurl:

0.2*40=8?

dude:

Yes Since that is a discount, subtract 8 from the original price Cost - discount = new cost 40-8

PuzzledGurl:

40-8=32 i did it in my head so it might be wrong

dude:

That is right :P

dude:

Okay now we need more 1 part it says there is a 5% tax after the discount

dude:

You want to repeat a similar pattern by finding what 5% of 32 is

PuzzledGurl:

That is the hard part okay so... how???? omg i am dumb

dude:

What is 5% as a decimal?

PuzzledGurl:

0.05 okay so do i basically do what i did

dude:

Multiply them now \(0.05\times32\)

PuzzledGurl:

0.5x32=16

dude:

Watch out with decimals It is 0.05 times 32, not 0.5 times 32

PuzzledGurl:

shoot

PuzzledGurl:

0.8

dude:

Might want to re-check that \(0.05\times 32\) should be 1.6

PuzzledGurl:

okay hold on i am so stupid i cant even calculate this OML Thanks

PuzzledGurl:

next

dude:

It's fine, everyone makes typing mistakes just try to make sure you type it correctly once you put it in Since this is tax, it is added to the cost So do 32+1.6 =

PuzzledGurl:

32+1.6=33.6 right

dude:

Looks good, that is your answer

PuzzledGurl:

so the price is $33.60?

dude:

Yeah

PuzzledGurl:

Will I have to open a new question for you to help me solve the next one

dude:

Sure

PuzzledGurl:

But i cant figure out how to open another math one OML I AM STOOPID

dude:

Close this question and then you should be able to post another one

PuzzledGurl:

how shall i close it

PuzzledGurl:

i figured it out